OpenStudy (anonymous):

What is the radius of a circle with circumference 22m? Instructions: Solve Each Problem. Round to the nearest tenth, If necessary. Use 3.14

OpenStudy (mathstudent55):

The formula for the circumference of a circle is: C = 2(pi)r You have C, and you are looking for r. pi = 3.14 22 m = 2(pi)r 2(pi)r = 22 m 2(3.14)r = 22 m 6.28r = 22 m Divide both sides by 6.28 r = (22/6.28) m Use your calculator to do the division.

OpenStudy (anonymous):

Thank you i got 3.503184713

OpenStudy (mathstudent55):

That's correct, but the instructions say round to nearest tenth.

OpenStudy (anonymous):

Yup i know

OpenStudy (anonymous):

Can you help me with my next question?

OpenStudy (mathstudent55):

New post pls